OpenStudy (anonymous):

what is the factorization of -x2+2x+48

OpenStudy (anonymous):

What are two numbers tat multiply to -48 and add up to 2?

OpenStudy (anonymous):

that*

OpenStudy (anonymous):

8,6

OpenStudy (anonymous):

is it (-x+6)(x+6)

OpenStudy (anonymous):

So you'll need something like:\[(-x \pm 8)(x\pm 6).\]You just have to figure out if its a plus or minus that goes in each one.

OpenStudy (uri):

-x2+2x+48 -x^2+8x-6x+48 -x(x+8)-6(x+8) (-x-6)(x+8)

OpenStudy (anonymous):

Since you want a positive 2x, I suggest trying:\[(-x+8)(x+6)\]

OpenStudy (campbell_st):

@uri... your epansion gives the coefficient of x as -14

OpenStudy (uri):

How? +8-6=2

OpenStudy (anonymous):

when you factor say \[-x^2+8x\]it becomes:\[-x(x-8)\]

OpenStudy (campbell_st):

@uri (-x -6)(x +8) is your solution...

OpenStudy (anonymous):

you have -x(x+8). Same with the right side, your missing a negative.

OpenStudy (uri):

Oh yeah..blurry sorry typo.. It is.. (-x-6)(x-8)

OpenStudy (uri):

2 a.m...I need my glasses xD

OpenStudy (campbell_st):

most people are happy factoring when the leading term is positive. So a simple method is the take -1 as a common factor... \[-1(x^2 -2x - 48)\] now look at the quadratic you need to factor... and use the information above.

OpenStudy (mathstudent55):

I suggest you first factor out -1: \(-x^2+2x+48\) \( = -1(x^2- 2x-48) \) Now you need 2 fators of -48 that add to -2: -8 and 6 \(= -1(x -8)(x + 6) \) Now if you'd like, you can multiply the -1 back in: \(= (-x + 8)(x + 6) \)